Dan receives $375 per week for selling new and brand new and used cars at a car dealership in oak lawn, illinois. In addition, he receives 5% of the profit on any sales that he generates. Write a linear equation that relates Dan's weekly salary S when he has sales that generate profit of x dollars.
Answer by mananth(16949)

5% is the income on sales on profit made onsales.
actual commission received = 5/100 * x
=x/20
S=375+x/20